### v1

# Dynamical symmetry breaking in Georgi-Glashow chiral-gauge theories

Abstract
We investigate dynamical symmetry breaking in a class of chiral gauge theories containing the Georgi-Glashow model. These theories feature a gauge sector and two fermion species that transform in the two-index antisymmetric and antifundamental representations with different multiplicities. Using the effective action formalism and the functional renormalization group, we derive the flow of four-fermion interactions that encode their resonant structure and information about bound-state formation. Generalizing the theories to multiple generations, we make contact with the loss of asymptotic freedom and dissect the boundary of a conjectured conformal window. Our results show that, while most of the theory space displays a dominant color-breaking condensate, there exists a strongly coupled regime where the lowest-laying mechanisms fail and more intricate dynamics are expected to arise. This analysis provides a first step toward the infrared behavior of chiral gauge theories with functional methods.
preprint: RIKEN-iTHEMS-Report-25
I Introduction
One of the outstanding challenges in Quantum Field Theory (QFT) and particle physics is to systematically understand the infrared (IR) phases of gauge theories in (3+1)(3+1)( 3 + 1 ) spacetime dimensions. A wide array of approaches, both exact and approximate, have been developed to probe the phases of such theories. Constraints have been developed from tools as the ’t Hooft anomaly matching conditions ’t Hooft (1980) or via connections to the supersymmetric limit Seiberg and Witten (1994a, b); Seiberg (1994, 1995). For vector-like gauge theories, the fermion path integral measure can be made real and positive by appropriately choosing fermion mass terms, thereby permitting the application of additional criteria such as the persistent mass condition [^1] See e.g. Ciambriello et al. (2024, 2025) for recent studies on the use of persistent mass conditions together with ’t Hooft anomaly matching conditions in a class of QCD-like theories, where it is proven that chiral symmetry must be spontaneously broken once the theory flows to an IR-free fixed point with full screening of color charges. Vafa and Witten (1984). Moreover, major insights into physical correlations in the strongly coupled regime have been obtained from first-principle computations such as Monte Carlo lattice simulations (see e.g. Wilson (1974); Kogut (1983); Greensite (2003); Bogolubsky et al. (2009); Bowman et al. (2004); Aoki et al. (2009)) and functional methods (see e.g. Alkofer and von Smekal (2001); Fischer (2006); Pawlowski et al. (2004); Braun et al. (2010); Cyrol et al. (2016); Aguilar et al. (2023)).
Although most efforts have focused on gauge theories with vector-like fermions, due to its relevance to Quantum Chromodynamics (QCD) Gross et al. (2023), the study of chiral gauge theories remains of significant theoretical and phenomenological interest. On the one hand, a deeper understanding of chiral gauge theories is essential, as the Standard Model (SM) of particle physics is constructed based on a chiral matter content. However, strongly coupled chiral gauge theories Raby et al. (1980); Dimopoulos et al. (1980); Bars and Yankielowicz (1981) may potentially be relevant for addressing outstanding questions beyond the SM. However, our limited understanding of their strongly coupled behavior represents a major obstacle to assessing their phenomenological viability. Accordingly, it is of paramount importance to advance our knowledge of the dynamics governing these theories.
By construction, no gauge-invariant fermion bilinear mass terms exist in chiral-gauge theories, which renders the persistent mass conditions inapplicable. Furthermore, the lattice regularization of chiral fermions remains a subtle and technically challenging task due to the Nielsen–Ninomiya theorem Nielsen and Ninomiya (1981a, b, c). These challenges have motivated a long-standing effort to develop robust theoretical tools capable of systematically probing the IR phases of chiral gauge theories. Recent progress has been made by exploiting ’t Hooft anomaly matching conditions associated with generalized global symmetries; for reviews, see e.g. Bolognesi et al. (2022, 2023); Konishi et al. (2024, 2024). Despite these advances, our understanding of chiral gauge dynamics remains incomplete, as either analytical or numerical methods are still largely unavailable.
The IR phases of such theories are expected to depend crucially on the fermion content Sannino (2009); Poppitz and Unsal (2009); Bolognesi and Konishi (2019); Bolognesi et al. (2020, 2021, 2022, 2023, 2024). Some models may flow to an interacting IR fixed point, while others are conjectured to exhibit non-trivial Higgsing dynamics with condensate formation, leading to rich phenomena such as color-flavor locking, tumbling, and dynamical Abelianization. The standard lore is that one can smoothly interpolate between the Higgs and confinement regimes without encountering a phase transition Fradkin and Shenker (1979); Dimopoulos et al. (1980).
In this paper, we study a class of anomaly-free and purely chiral gauge theories of the Georgi-Glashow (GG) type using the functional renormalization group (fRG) approach Wetterich (1993a); Dupuis et al. (2021). Due to its non-perturbative nature, the fRG is particularly suitable for investigating the strongly coupled dynamics of these theories. In this first work, we focus on the dynamical symmetry breaking (dSB) mechanisms which give rise to the highest condensate in the theory following in spirit previous studies in QCD-like theories Miransky (1999); Miransky and Yamawaki (1997); Gies and Jaeckel (2006); Braun and Gies (2006, 2010); Goertz et al. (2024); Braun et al. (2011). While here we focus on the start of dynamics, the present approach has proven its quantitative precision to describe dynamical chiral symmetry breaking and color confinement. Remarkably, it is shown in Cyrol et al. (2018); Mitter et al. (2015); Ihssen et al. (2024); Goertz et al. (2024); Fu et al. (2020) that the fRG approach can achieve quantitative precision with lattice results where they exist. Consequently, no conceptual complication is expected in the present approach.
In this work, we derive the RG flows for the dressings of the four-fermion operators which provide us with knowledge on the channel’s resonant structure and the critical coupling necessary to trigger a dynamical scale generation. To the best of our knowledge, this class of multispecies gauge–fermion theories, where each fermion transforms under distinct, nonconjugate representations of the gauge group, has not been previously studied using functional methods. This provides us with unprecedented information on the properties of the theory at the start of non-perturbative dynamics as well as allows us to chart theory space where various regions with distinct dynamical patterns can be identified, including a conjectured conformal limit.
The remainder of the paper is organized as follows. In Section II we introduce the Georgi-Glashow model and its generalized form, which allows us to make contact to the loss of asymptotic freedom and a quantum scale-invariant regime. In Section III we introduce the effective action approach, the methodology for diagnosing dSB, and the fRG approach. In Section IV we discuss the results of the flow equation, concerning the resonant structure of the different channels and the phases in theory space, including the quantum scale invariant limit. Before concluding, we speculate about the IR fate of the theories and discuss the connection to other studies. Technical details concerning the basis construction and the flow equations as well as single examples can be found in the Appendices.
II The chiral theories
We are interested in purely chiral theories and in this work we focus on a general class consisting of $SU(N_{c})$ gauge theories ($N_{c}\geq 5$) with left-handed Weyl fermions in the respective complex representations,
[…]
where the fermion multiplicities are related in order to cancel the $[SU(N_{c})]^{3}$ triangle gauge anomaly. When $N_{c}=5$, the fermion content coincides with the model originally proposed by Georgi and Glashow in the context of grand unification of the SM gauge couplings Georgi and Glashow (1974). Due to this, models of this type with $N_{\textrm{gen}}=1$ are later referred to as chiral-gauge theories of the Georgi-Glashow type Raby et al. (1980); Dimopoulos et al. (1980), arguably the simplest purely chiral models.
Motivated by the structural similarity to the SM, throughout the rest of the paper we refer to $N_{\textrm{gen}}$ as the number of generations or copies of the minimal GG theory. Within each generation, the model contains one Weyl fermion in the two-index antisymmetric representation, $\chi$, and accordingly, $(N_{c}-4)$ Weyl fermions in the anti-fundamental representation, $\psi$. The GG model with $N_{\textrm{gen}}$ generations of chiral fermions has the global flavor symmetry group
$G_{f}=SU\left(N_{\textrm{gen}}\left(N_{c}-4\right)\right)\times SU\left(N_{\textrm{gen}}\right)\times U\left(1\right)\;,$ (2)
up to some central group elements acting trivially on the chiral fermions; see Table 1 for the various fermion charges. Notice that $G_{f}$ reduces to a single $U(1)$ when $N_{c}=5$ and $N_{\textrm{gen}}=1$, otherwise there are always non-Abelian factors.
$SU(N_{c})$ $\,\,SU\left(N_{\textrm{gen}}(N_{c}-4)\right)\,\,$ $\,\,SU\left(N_{\textrm{gen}}\right)\,\,$ $U(1)$
[…]
Table 1: Fermion content of the generalized GG model with $N_{\textrm{gen}}$ generations, where the fermion quantum numbers are specified under the $SU(N_{c})$ gauge group and the flavor symmetry group defined in Eq. (2). Here, the $U(1)$ factor is free from the $U(1)[SU(N_{c})]^{2}$ Adler-Bell-Jackiw anomaly.
Theories with $N_{\textrm{gen}}<N_{\textrm{gen}}^{\rm AF}=(11N_{c})/(2N_{c}-6)$ have a well-defined asymptotically free ultraviolet (UV) limit. When $N_{\textrm{gen}}$ is slightly below $N_{\textrm{gen}}^{\rm AF}$, it is believed that theories flow into an interacting IR fixed point, analogous in spirit to the Caswell-Banks-Zaks Caswell (1974); Banks and Zaks (1982) solutions in QCD-like theories. However, unlike the latter, determining the existence of the fixed point in the GG-type models is tricky. In particular, a perturbative analysis of the existence of a fixed point is not valid, since the zeros of the $\beta$ function cannot be parametrically under control [^2] It may still be numerically small, but this is not sufficient to prove the existence of an interacting fixed point in the perturbative regime., contrary to QCD-like theories in the Veneziano-Witten limit. A simple estimation using the leading two coefficients of the $\beta$ function suggests that all the interacting fixed points, if they exist at all, are non-perturbative, even when the gauge coupling can be numerically small; see Bolognesi et al. (2023) for a review emphasizing this point. For sufficiently small $N_{\textrm{gen}}$, fixed-point solutions are not found and the gauge coupling drives the appearance of so-far unknown dynamics which we target in this work.
We note that while the strongly coupled regime of the one-generation model was studied in Raby et al. (1980); Dimopoulos et al. (1980) long ago, the multi-generation GG models, which we refer to as the generalized, have received very little attention; for instance, see only Bolognesi and Konishi (2019); Bai and Stolarski (2022) for some recent studies.
III Effective action and the functional approach
In the present work, we employ the effective action formalism which encodes all physical information in the one-particle irreducible generating functional, from which correlation functions can be extracted. These fully-dressed correlators or $n$-point functions capture the relevant dynamics of the theory, including known phenomena like confinement and dynamical chiral symmetry breaking, as observed in QCD, as well as potential new mechanisms not yet identified. A particularly important feature in gauge–fermion quantum field theories is the dynamical formation of a condensate, which leads to the generation of an explicit mass scale. In the present classically scale-invariant theories, this also encodes the breaking of quantum scale invariance, which may otherwise be given in the presence of an IR fixed point, as in the CBZ scenario.
The dynamical emergence of fermion condensates has a long history in the context of effective actions, particularly in condensed matter and nuclear physics Nambu and Jona-Lasinio (1961); Braun (2012). This mechanism is encoded in the fermionic $n$-point functions and can commonly be diagnosed in a divergence of the dressings. Through a Hubbard–Stratonovich transformation Hubbard (1959); Stratonovich (1957), this divergence is shown to be mathematically equivalent to the appearance of a condensate in the bosonized formulation of the theory. In that setting, the formation of a condensate corresponds to a change in the curvature of the bosonic effective potential around the origin, specifically, a change in the sign of the boson mass term. Thus, the transition from a symmetric to a broken phase is indicated by a vanishing boson mass (i.e., an infinite correlation length), or equivalently, by a divergence in the corresponding four-fermion coupling. See, e.g., Miransky and Yamawaki (1989); Miransky (1994); Gies and Jaeckel (2006) for illustrative examples.
In this work, we follow these lines to analyze the appearance of condensates in generalized GG models. To account for quantum corrections and derive the $n$-point functions of the effective action, we employ the functional renormalization group Wetterich (1993a). In this non-perturbative approach, momentum modes in the path integral measure are integrated out progressively with the help of an IR regulator function $R_{k}$, which suppresses fluctuations below the cutoff scale $k$. This procedure defines the effective average action $\Gamma_{k}$ Wetterich (1991, 1993b), which incorporates quantum fluctuations down to the momentum scale $k$ and recovers the full effective action in the limit $k\to 0$. Central to this approach is the flow or Wetterich equation with which we implement quantum corrections in a non-perturbative manner,
$\partial_{t}\Gamma_{k}\left[\Phi\right]=\frac{1}{2}\,{\rm STr}\left[\left(\Gamma_{k}^{(2)}\left[\Phi\right]+R_{k}\right)^{-1}\cdot\partial_{t}R_{k}\right]\,.$ (3)
Here, $\Phi$ denotes the superfield containing all fields in the theory, $\partial_{t}\equiv k\,\partial_{k}$ is the RG-time derivative, and $\Gamma_{k}^{(2)}\left[\Phi\right]$ is the second functional derivative of $\Gamma_{k}\left[\Phi\right]$, i.e. the full inverse propagator Hessian. The supertrace $\mathrm{STr}$ includes a trace over momentum space as well as all internal (gauge and global) indices, with a minus sign for fermionic degrees of freedom.
Exactly solving the flow equation (3) is generally not feasible, and practical applications rely on truncations of the effective average action. These consist of considering a finite set of operators consistent with the symmetries of the theory, each with scale-dependent dressings. In this work, we consider the minimal truncation including up to four-fermion operators which permit diagnosing spontaneous symmetry breaking. The regulator function $R_{k}$ is not unique and can be chosen to optimize convergence and numerical stability. We employ the Litim regulator Litim (2001) which provides analytic handle on the flow equations. We refer technical details on the flow equation and truncation to Appendix B.
Functional approaches are widely used to study dSB, particularly in QCD and its phase structure, which are close in spirit to the theories studied here. Advanced fRG computations incorporate explicit momentum dependence in higher $n$-point functions, both in vacuum and at finite temperature or chemical potential, see e.g. Fu et al. (2023, 2024, 2025); Braun et al. (2018, 2020); Ihssen et al. (2024). Moreover, the fRG framework allows for scale-dependent field transformations along the flow, enabling the description of dynamically emerging effective degrees of freedom Gies and Wetterich (2004); Fukushima et al. (2021); Pawlowski (2007). The emergent composites (or dynamical hadronisation) approach facilitates the inclusion of higher-dimensional fermionic operators via bosonic interactions $\phi^{n}\propto\left(\bar{\psi}\,{\cal T}\psi\right)^{n}$, thus extending the analysis beyond the singularity of the four-fermion coupling that signals dSB and providing smooth access to the chirally broken regime. This formulation offers both technical and conceptual advantages for studying strongly coupled gauge–fermion systems.
In this first work, we restrict ourselves to a qualitative analysis and leave the full bosonized treatment to future studies. Our goal is to identify which condensate forms at the highest scales, which is the first step towards the full IR dynamics. For this, it suffices to employ a minimal Fierz-complete truncation of four-fermion operators Gies and Jaeckel (2006); Braun and Gies (2007, 2006); Braun et al. (2011).
The effective average action for generalized GG gauge-fermion theories has the form,
$\Gamma_{k}\left[A_{\mu},\bar{c},c,\bar{\psi},\psi,\bar{\chi},\chi\right]=$ $\,\Gamma_{{\rm gauge},k}\left[A_{\mu},\bar{c},c\right]$ (4)
$\hskip-99.58464pt+\Gamma_{{\rm Dirac},k}\left[A_{\mu},\bar{\psi},\psi,\bar{\chi},\chi\right]+\Gamma_{{\rm 4F},k}\left[\bar{\psi},\psi,\bar{\chi},\chi\right]+\ldots\,,$
where $\ldots$ account for the remaining higher-dimensional operators which are present in the full effective action but are not taken into account in the minimal truncation here. The first two terms in (4) stand for the pure gauge and Dirac parts including up to dimension four operators (e.g. dispersions and interaction terms). As mentioned, the information on dSB and bound state formation is encoded in the purely fermionic operators which in (4) constitute
$\Gamma_{{\rm 4F},k}[\bar{\psi},\psi,\bar{\chi},\chi]=$ (5)
$\quad-\int_{x}Z_{\psi}^{2}\,\sum_{i=1}^{2}\lambda_{i}{\cal O}_{i}+Z_{\chi}^{2}\sum_{i=3}^{5}\lambda_{i}{\cal O}_{i}+Z_{\psi}Z_{\chi}\sum_{i=6}^{7}\lambda_{i}{\cal O}_{i}\,,$
and include all possible four-fermion operators forming a Fierz-complete basis which we construct with the Young Tensor Method Li et al. (2021a, b, 2022). In (5), $Z_{\psi}$ and $Z_{\chi}$ are the fermionic field wave functions and $\lambda_{i}=k^{-2}\bar{\lambda}_{i}$ are the dimensionless form of the four-fermion dressings for the set of operators
${\cal O}_{1}$ $=\big{(}\psi^{\dagger}\bar{\sigma}^{\mu}\psi\big{)}\big{(}\psi^{\dagger}\bar{\sigma}^{\mu}\psi\big{)},$ (6a)
${\cal O}_{2}$ $=\big{(}{\psi^{\dagger}}^{f_{1}}\bar{\sigma}^{\mu}\psi_{f_{2}}\big{)}\big{(}{\psi^{\dagger}}^{f_{2}}\bar{\sigma}^{\mu}\psi_{f_{1}}\big{)},$ (6b)
${\cal O}_{3}$ $=\big{(}{\chi^{\dagger}}^{f_{1}}\bar{\sigma}^{\mu}\chi_{f_{2}}\big{)}\big{(}{\chi^{\dagger}}^{f_{2}}\bar{\sigma}^{\mu}\chi_{f_{1}}\big{)},$ (6c)
${\cal O}_{4}$ $=\big{(}{\chi^{\dagger}}\bar{\sigma}^{\mu}\chi\big{)}\big{(}{\chi^{\dagger}}\bar{\sigma}^{\mu}\chi\big{)},$ (6d)
${\cal O}_{5}$ $=\big{(}{\chi^{\dagger}}\bar{\sigma}^{\mu}\,T_{\rm anti}\,\chi\big{)}\big{(}{\chi^{\dagger}}\bar{\sigma}^{\mu}\,T_{\rm anti}\,\chi\big{)},$ (6e)
${\cal O}_{6}$ $=\big{(}\psi^{\dagger}\bar{\sigma}^{\mu}\psi\big{)}\big{(}{\chi^{\dagger}}\bar{\sigma}^{\mu}\chi\big{)},$ (6f)
${\cal O}_{7}$ $=\big{(}\psi^{\dagger}\bar{\sigma}^{\mu}\,T_{\rm a-fund}\,\psi\big{)}\big{(}{\chi^{\dagger}}\bar{\sigma}^{\mu}\,T_{\rm anti}\,\chi\big{)}\,.$ (6g)
Here, we adopted the convention where fermionic color and the flavor ($f_{\ldots}$) indices are assumed contracted within each parenthesis unless explicitly displayed. $T_{\rm anti}$ and $T_{\rm a-fund}$ stand for the color group generators in the anti-symmetric and anti-fundamental representations, respectively. To construct this Fierz-complete basis we have imposed invariance under the Lorentz group, the gauge group $SU\left(N_{c}\right)$ and the whole flavor group $SU\left(N_{\textrm{gen}}\left(N_{c}-4\right)\right)\times SU\left(N_{\textrm{gen}}\right)\times U\left(1\right)$, while taking into account the effect of repeated fields and Fierz redundancy, see Appendix A for details on the construction.
The basis in (6) includes all fermionic self-interaction terms at canonical dimension six, with the exception of the operator which appears only in the theory with $N_{c}=5$ and $N_{\rm gen}=1$:
${\cal O}_{8}=\epsilon_{i_{1}i_{2}i_{3}i_{4}i_{5}}\left(\chi^{i_{1}i_{2}}\chi^{i_{3}i_{4}}\right)\left(\chi^{i_{5}j}\psi_{j}\right),$ (7)
where $\epsilon_{i_{\dots},\ldots}$ is the Levi-Civita tensor for the color group. This operator corresponds to the so-called ’t Hooft vertex, where the number of fermions is determined by the Dynkin indices of the fermions under the gauge group, which is associated with the axial anomaly and counts the number of fermion zero modes in the instanton background with unit topological charge. For theories with generic values of $N_{\textrm{gen}}$ and $N_{c}$, the operator corresponding to the ’t Hooft vertex involves $(N_{c}-4)N_{\textrm{gen}}$ Weyl fermions of $\psi$ and $(N_{c}-2)N_{\textrm{gen}}$ Weyl fermions of $\chi$, hence the operator has canonical mass dimension $\left[{\cal O}_{8}\right]=3(N_{c}-3)N_{\textrm{gen}}$.
Since we focus on the large $N_{\textrm{gen}}$ limit, we omit this operator in the present analysis as becomes increasingly canonically irrelevant. We further note that for the study of the single-generation case, the present operator basis further simplifies as ${\cal O}_{1}\equiv{\cal O}_{2}$ and ${\cal O}_{3}\equiv{\cal O}_{4}$.
IV Patterns of symmetry breaking
The evolution of four-fermion dressings is dictated by their respective RG flows, which are triggered purely by the gauge dynamics. While these flows have been extensively studied in QCD-like theories (see e.g. Goertz et al. (2024); Fu et al. (2023, 2024, 2025); Braun (2012)), as we will see, the situation in chiral-gauge theories with several species is qualitatively different and remains rather unexplored. We commence this Section by examining the general structure of the four-fermion RG flow in gauge–fermion systems and the underlying driving mechanism. We then discuss the resonant structure in this class of theories and provide more details about particular theories in Appendix C.
The RG flow of a general four-fermion coupling has the form,
$\partial_{t}\bar{\lambda}_{i}\propto$  $\,2\,\bar{\lambda}_{i}+\boldsymbol{c}_{{\rm A},i}\cdot\alpha_{g}^{2}+\boldsymbol{c}_{{\rm B},ij}\cdot\alpha_{g}\,\bar{\lambda}_{j}$
$\hskip 85.35826pt+\boldsymbol{c}_{{\rm C},ijk}\cdot\bar{\lambda}_{j}\bar{\lambda}_{k}+\ldots\,,$ (8)
where $\alpha_{g}=g^{2}/4\pi$ and we have only explicitly shown the lowest-order contributions. As the theory flows away from the UV Gaussian fixed point (the free theory), quantum fluctuations generate all kind of operators allowed by the symmetries via gauge interactions. More precisely, the four-fermion operators are triggered through gauge-boson-mediated box diagrams encoded in $\boldsymbol{c}_{{\rm A},i}$ in (8). Once generated, these feed into their own flows via the terms proportional to $\boldsymbol{c}_{{\rm B},ij}$, and the $\boldsymbol{c}_{{\rm C},ijk}$ terms control the self-interaction structure. The indices $j,k$ here, range over all channels including $i$ that corresponds to the basis in (6). Since fermionic self-interactions are triggered by gauge fluctuations, they may be interpreted as higher-order gauge corrections, with the various dressings corresponding to distinct tensor channels. For further details, see Braun (2012); Goertz et al. (2024); Dupuis et al. (2021) and references therein. The coefficients encode the color and global symmetries of the theory and display a non-trivial dependence on $N_{c}$ and $N_{\textrm{gen}}$, resulting from tracing over the internal structure of the theory. Finally, the first term in (8) accounts for the canonical dimension of the operator.
In total, the self-interaction terms $\boldsymbol{c}_{{\rm C},iii}$ yield a parabola shape (see Figure 4 and the respective particular example in Appendix B) of the flow as a function of its respective dressing which, depending on the sign of the coefficient can yield an upright or inverted shape. Moreover, for small $\alpha_{g}$ the flow of the dressings displays an IR attractive fixed point which in the classical limit ($k\to\infty$) reduces to the Gaussian one, $\alpha_{g}^{*}=0$ and $\bar{\lambda}_{i}^{*}=0$.
As the gauge dynamics evolve towards the IR, the parabola shape is shifted upwards or downwards depending on the sign of the $\boldsymbol{c}_{{\rm A},i}$ coefficients, which read
[…]
Note that none of these depend on $N_{\textrm{gen}}$, but only on $N_{c}$ (see Figure 2).
In order for a singularity to develop, the gauge dynamics must drive the disappearance of the IR fixed point, i.e., shift the RG flow parabola such that its zeros merge and vanish. This requires that the leading gauge corrections ($\boldsymbol{c}_{{\rm A},i}$) and the higher-order self-interactions ($\boldsymbol{c}_{{\rm C},iii}$) act coherently, namely by having the same sign, to induce a fixed-point annihilation. While this alignment offers a qualitative criterion for identifying a resonant structure, it relies on a simplified scenario where inter-channel contributions are subleading and the flow is dominated by the leading gauge terms. More intricate dynamics arise when off-diagonal terms become relevant, such as $\boldsymbol{c}_{{\rm B},ij}$ with $i\neq j$ or $\boldsymbol{c}_{{\rm C},ijk}$ with $i\neq j,k$, which can significantly alter the flow and the mechanism of symmetry breaking. These cases will be encountered below.
Moreover, the presence of a resonant structure in a given channel is a necessary but not sufficient condition for a singularity to occur, as the gauge coupling must also reach a sufficient strength to cause the fixed-point merger. To quantify this, we define the critical coupling $\alpha^{\textrm{crit}}_{\textrm{SB}}$ as the minimal value of the gauge coupling $\alpha_{g}$ for which at least one of the four-fermion flows ceases to exhibit a root. Explicitly,
$\alpha^{\textrm{crit}}_{\textrm{SB}}=$ (10)
$\hskip-28.45274pt\inf\left\{\alpha_{g}\,\middle|\,\exists\,i:\,\,\left(\,\partial_{t}\bar{\lambda}_{i}\leq 0\quad\forall\,\bar{\lambda}_{j}\right)\,{\,\rm or\,}\left(\,\partial_{t}\bar{\lambda}_{i}\geq 0\quad\forall\,\bar{\lambda}_{j}\right)\right\}\,,$
which is equivalent to the maximum $\alpha_{g}$ at which all flows remain bounded at a fixed $k$. It follows from (10) that for $\alpha_{g}<\alpha^{\textrm{crit}}_{\textrm{SB}}$, all four-fermion couplings $\bar{\lambda}_{i}$ flow to IR fixed points and for $\alpha_{g}>\alpha^{\textrm{crit}}_{\textrm{SB}}$, at least one coupling ceases to have a fixed point and flows towards a singularity, signaling the onset of dSB. The dominant channel is then identified as the one whose interaction term drives the flow away from the IR fixed point, thereby setting the dominant singularity of the system.
For the class of theories considered here, the diagonal coefficients $\boldsymbol{c}_{{\rm C},iii}$ read:
$\boldsymbol{c}_{{\rm C},\,iii}$ $=\frac{1}{8\pi^{2}}\bigg{\{}\;N_{c}N_{\textrm{gen}}\left(N_{c}-4\right)-1,\;N_{c}+N_{\textrm{gen}}\left(N_{c}-4\right),\;2\,(N_{c}-1)N_{c}+4\,N_{\textrm{gen}},\;\,\frac{(N_{c}-1)N_{c}N_{\textrm{gen}}}{2}-1,$ (11)
$\hskip 213.39566pt\,\frac{N_{c}\,(N_{\textrm{gen}}+3)}{2}+\frac{8}{N_{c}}-\,N_{\textrm{gen}}+1,\;-\frac{3}{2},\;\frac{N_{c}}{4}+\frac{3}{N_{c}}+\frac{3}{4}\bigg{\}}\,.$
Although these coefficients retain a constant sign for $N_{c}>5$, the associated $\boldsymbol{c}_{{\rm A},i}$ undergo a sign change. Recalling the resonance condition, this observation indicates that certain channels may become resonant in specific regions of theory space but not in others, thereby anticipating distinct channel dominance and symmetry-breaking patterns across different domains.
Refer to caption     Refer to caption
Figure 1: The three-dimensional plot on the left shows, as a colored surface, the critical strength of the gauge coupling ($\alpha^{\textrm{crit}}_{\textrm{SB}}$) required to trigger the dynamical generation of a scale in the space of generalized Georgi–Glashow theories, spanned by the number of colors ($N_{c}$) and generations ($N_{\rm gen}$). The right-hand plot displays the same surface projected onto the plane and different regimes are labeled. The solid black line marks the boundary of asymptotic freedom and the dashed, dashed-dotted and dotted pink lines indicate the theories where the $\alpha^{\textrm{crit}}_{\textrm{SB}}$ computed coincides with the IR fixed points from the two- (dashed), three- (dashed-dotted) and four-loop (dotted) $\overline{\text{MS}}$ beta functions, respectively. The shaded region above these curves shows the conjectured conformal window.
To explore the $N_{c}-N_{\textrm{gen}}$ landscape, we compute $\alpha^{\textrm{crit}}_{\textrm{SB}}$ by numerically solving the system of flow equations in Appendix B. This analysis reveals the existence of distinct regions in theory space (see Figure 1):
  • $\bullet$
    Region I: For $N_{c}\lesssim 11$, only the channels ${\cal O}_{5}$ and ${\cal O}_{7}$ exhibit the discussed naive resonant structure. Among these, ${\cal O}_{5}$ shows a clear dominance which means that its evolution is not affected by the other channels and that is the one which develops the singularity at the highest scale. Furthermore, the $\alpha^{\textrm{crit}}_{\textrm{SB}}$ condition is satisfied for the respective dressing first. The other channels have a negligible role in the dynamics and dropping them does not have an impact on the highest scale of symmetry breaking nor on the qualitative nature of the condensate. Moreover, as illustrated in Figure 1, $\alpha^{\textrm{crit}}_{\textrm{SB}}$ remains essentially unchanged up to the vicinity of loss of asymptotic freedom. This result suggests that the leading condensate is $\langle\chi\chi\rangle$. Although this channel becomes singular at the highest scale and drives the initial symmetry-breaking dynamics, a more detailed IR analysis is required to determine the gauge representation of the condensate and the structure of its vacuum expectation value. Beyond this first scale, the next relevant interaction, ${\cal O}_{7}$, may give rise to a $\langle\psi\chi\rangle$ condensate. Interestingly, its resonant structure disappears around $N_{c}\sim 7$, potentially indicating the presence of subregions with qualitatively different IR behavior. Furthermore, the absence of resonance in the ${\cal O}_{1}$${\cal O}_{4}$ channels suggests that $\langle\psi\psi\rangle$ condensates are unlikely to form. Nonetheless, to make definitive statements about the subsequent dynamical scales, a careful treatment of the decoupling of massive modes, both fermionic and gauge, is required.
As $N_{c}$ increases, the resonant structure in both ${\cal O}_{5}$ and ${\cal O}_{7}$ is lost due to the sign change in the associated $\boldsymbol{c}_{{\rm A},i}$’s. Around $N_{c}\sim 11$, the simple mechanism ceases to apply in any channel, yet a singularity still develops via a more involved process. For several generations ($N_{\rm gen}\gtrsim 5$ and $N_{c}\gtrsim 11$), off-diagonal contributions $\boldsymbol{c}_{{\rm B},5j}$ and $\boldsymbol{c}_{{\rm C},5jk}$ restore the resonant flow in $\bar{\lambda}_{5}$, effectively leading to the scenario in Region I but with a mild increase in $\alpha^{\textrm{crit}}_{\textrm{SB}}$. The fixed-point merger is now driven by higher-order terms (e.g. $\alpha_{g}\bar{\lambda}_{j}$ and $\bar{\lambda}_{i}\bar{\lambda}_{j}$) rather than the original $g^{4}$. This limit is more non-perturbative, and the key contributions arise from channels such as ${\cal O}_{3}$ and ${\cal O}_{4}$ feeding into ${\cal O}_{5}$. However, as will be discussed below, most of this region likely falls within the conformal window, as the gauge coupling required to trigger symmetry breaking is comparable to or exceeds the values estimated for the IR fixed points.
However, for large $N_{c}\gtrsim 11$ and small $N_{\textrm{gen}}$, the off-diagonal contributions in the flow are too weak to induce a resonance in the previously dominant ${\cal O}_{5}$ channel. This transition reflects a change in the quantum realization of the gauge and global symmetries. Here we identify a further regime, whose precise boundaries depend sensitively on both $N_{c}$ and $N_{\textrm{gen}}$:
  • $\bullet$
    Region II: Here, all $\boldsymbol{c}_{{\rm A},i}$ are negative, and the four-fermion couplings flow to large negative values without forming a resonant structure in any single channel. Eventually, all $\bar{\lambda}_{i}\sim\mathcal{O}(1)$, and several couplings diverge simultaneously. In this regime, the system is highly entangled, and no clear dominant channel can be distinguished. We attempted to diagonalize the flow equations to define a basis $\tilde{\lambda}_{i}$ such that each $\tilde{\lambda}_{i}$ is independent of $\tilde{\lambda}_{j}$ for $j\neq i$, but found this not feasible. This regime features a novel mechanism for scale generation driven by high-order interactions, potentially involving exotic higher-dimensional condensates. We note that here the use of the flow equation instead of a perturbative method is particularly beneficial to throw some light into the non-perturbative dSB mechanism. A further alternative scenario in this region concerns the appearance of a gauge-boson mass gap which screens all gauge-fermion interactions. This could prevent the reach of $\alpha^{\textrm{crit}}_{\textrm{SB}}$ and hence the appearance of any condensate, leading to an scenario with color confinement and no global symmetry breaking. The present framework allows for such an assessment Cyrol et al. (2016); Goertz et al. (2024) which will be explored in a future work.
In summary, we found a rich landscape of global symmetry-breaking patterns, which we summarize in Figure 1. The emergence of singularities and resonant channels depends sensitively on $N_{c}$, $N_{\rm gen}$, and the interplay of tensor structures. While bosonization clarifies the nature of the condensates, higher-dimensional operators may be needed to fully capture the dynamics in certain regimes. These results call for future work extending to larger $N_{c}$ and exploring exotic IR phenomena.
A further constraint remains to be addressed. Gauge–fermion theories are known to exhibit a conformal regime in the many-fermion limit, close to the boundary of asymptotic freedom. In this regime, a perturbative IR fixed point emerges, quantum scale invariance is realized, and no dynamical scale is generated.
By comparing $\alpha^{\textrm{crit}}_{\textrm{SB}}$ with the IR fixed-point values of the gauge beta function, one can estimate the lower boundary of the conformal window, following the approach used for QCD-like systems in Goertz et al. (2024); Gies and Jaeckel (2006). In fact, if the fixed point of the gauge beta function is below the derived $\alpha^{\textrm{crit}}_{\textrm{SB}}$, then the gauge dynamics can never trigger spontaneous symmetry breaking, and eventually flows to an IR-interacting fixed point. In Figure 1, we display the boundaries obtained with the IR fixed point of the two- (dashed), three- (dot-dashed), and four-loop (dotted) $\overline{\text{MS}}$ beta function. The four-loop coefficient, computed for the first time for this class of theories, is presented in Appendix D, along with further details relevant to this regime.
Given the recent interest in these models, particularly for $N_{\textrm{gen}}=1$ Bolognesi et al. (2024, 2020, 2021, 2022); Bolognesi and Konishi (2019); Csáki et al. (2021); Tong (2022); Razamat and Tong (2021); Cacciapaglia et al. (2019), we briefly make contact with the conjectured scenarios. For $N_{c}>5$, the most-attractive-channel arguments Raby et al. (1980); Eichten and Feinberg (1982) and the proposals Bolognesi et al. (2024, 2020, 2021, 2022); Bolognesi and Konishi (2019); Csáki et al. (2021) highlight a first potential $\langle\psi\chi\rangle$ condensate. The latter could break $SU(N_{c})\times SU(N_{c}-4)\rightarrow SU(4)_{c}\times SU(N_{c}-4)_{cf}$ via color-flavor locking, consistent with discrete ’t Hooft anomaly matching conditions. This is in disagreement with our first found condensate $\langle\chi\chi\rangle$, however we identified some clues that a $\langle\psi\chi\rangle$ condensate might also appears at lower energies. Moreover, it is possible that the first $\langle\chi\chi\rangle$ condensate also breaks the gauge group to $SU(4)$, leaving the low-energy theory identical. Finally, contrary to what was found in Csáki et al. (2021), we do not find evidence for the emergence of a $\langle\psi\psi\rangle$ condensate and the respective channels do not show any signature to fulfill the necessary resonant structure.
V Conclusions
In this work, we studied dynamical symmetry breaking in chiral-gauge theories, focusing on the emergence of fermion condensates that spontaneously break global and gauge symmetries, and consequently quantum scale invariance. Specifically, we have analyzed the class of generalized Georgi–Glashow theories, which contains two species of fermions transforming under different color representations and extended it to arbitrary numbers of generations to make contact to the loss of asymptotic freedom and a conjectured conformal window.
We constructed the Fierz-complete basis of four-fermion operators and derived their flows using the functional renormalization group. These carry the information on the resonant structure of the different channels and easily permit to compute the critical coupling necessary to trigger the dynamical generation of a scale. We find that the phase structure of generalized GG theories is rich as various regions with distinct patterns of dSB. Such richness is a consequence of considering a theory with different fermion species rather than from the inherently chiral nature, as this provides strong constraints on the inducible operators. We note that, to the best of our knowledge, such type of multi-species gauge-fermion theories have not been studied before with approaches sensitive to the off-shell dynamics.
In Region I, where $N_{c}\lesssim 11$, the dominant tensor channel ${\cal O}_{5}$ becomes resonant, suggesting the formation of a $\langle\chi\chi\rangle$ condensate. However, for a larger number of colors in Region II, the critical gauge coupling increases substantially, and the fixed-point merger is triggered by off-diagonal and higher-order contributions, suggesting a strongly coupled regime with qualitatively different underlying dynamics of non-perturbative nature. Furthermore, we provide an estimate on the boundary of the conjectured conformal window in the many generation limit, close to the loss of asymptotic freedom. This is obtained by comparing the derived critical coupling for dynamics to high-loop fixed-point values from perturbative $\beta$-functions, which the four-loop we derive in this work.
Finally, we briefly comment on possible IR scenarios beyond the first condensate, where the interplay between different channels, the pattern of symmetry breaking and potential higher-order condensates remain open questions. A detailed resolution of these dynamics, including the emergence of color-flavor-locked phases or multi-fermion condensates, will require bosonized treatments or more elaborate truncations.
The present analysis opens the door to many future studies. In an upcoming companion paper, we will extend this work in the symmetric case, commonly known as the Bars-Yankielowicz Bars and Yankielowicz (1981) model, exploring whether similar or qualitatively new symmetry-breaking patterns arise. Furthermore, the IR dynamics of these models can be tackled from first-principles in the present framework by dynamical bosonizing the dominant channel and including the dynamical appearance of a gauge-boson mass gap. More broadly, the functional method can be applied to investigate the mechanism of symmetric mass generation in various spacetime dimensions, see e.g. Wang and You (2022) for a review. Once all the ’t Hooft anomalies are canceled, this mechanism can be used to generate a mass gap for chiral fermions without breaking the chiral symmetry. See e.g. Wang and Wen (2023, 2018); Zeng et al. (2022); Tong (2022) for the studies on the chiral fermion 3 - 4 - 5 - 0 model in 1+1 dimensions, and Razamat and Tong (2021); Tong (2022) for other related models in 3+1 dimensions.
Acknowledgements.
We would like to thank Jan M. Pawlowski for helpful comments and suggestions. APG is supported by the RIKEN Special Postdoctoral Researcher (SPDR) Program. SV is supported by the Fonds de la Recherche Scientifique de Belgique (FNRS) under the IISN convention 4.4517.08. We are grateful to the Mainz Institute for Theoretical Physics (MITP) of the DFG Cluster of Excellence PRISMA+ (Project ID 39083149) for its hospitality and support during the initial stages of this work. H-.L.L. is supported by the start-up funding of Sun Yat-Sen University under grant number 74130-12255013. The work of L.X.X. is partially supported by “Exotic High Energy Phenomenology” (X-HEP), a project funded by the European Union - Grant Agreement n.101039756. This project is also supported by the Munich Institute for Astro-, Particle and BioPhysics (MIAPbP) which is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC-2094 – 390783311.
Appendices
Appendix A Four-fermion operator construction
This Appendix is devoted to the systematic construction of the four-fermion operators in generalized GG theories. We demand that operators are not only invariant under the $SU(N_{c})$ gauge group, but also under the global flavor symmetry for each chiral fermion and the residual anomaly-free $U(1)$ symmetry as well. This allows only three operator types: ${\chi^{\dagger}}^{2}\chi^{2}$, ${\chi^{\dagger}}\chi{\psi^{\dagger}}\psi$ and ${\psi^{\dagger}}^{2}\psi^{2}$, except for the scenario $N_{c}=5$ with $N_{\textrm{gen}}=1$ in the GG model, where one additional operator type (and its complex conjugate) is allowed: $\psi\chi^{3}$ (see discussion in III.
For operator types ${\psi^{\dagger}}^{2}\psi^{2}$ and ${\chi^{\dagger}}^{2}\chi^{2}$, the repeated-fields effect needs to be taken into account the operator basis construction. The repeated fields are the ones with the same gauge and Lorentz quantum numbers, e.g. two $\psi$’s in the type ${\psi^{\dagger}}^{2}\psi^{2}$. Due to the spin-statistics, operators should be totally anti-symmetric (symmetric) regarding to swapping any two of the repeated fields if the corresponding fields are fermions (bosons). Therefore, the permutation symmetry of the combined tensor of Lorentz, gauge and flavor should be totally anti-symmetric in these two types of operators. For ${\psi^{\dagger}}^{2}\psi^{2}$, two independent invariant tensors can be used to contract the flavor indices:
$T^{[f]}_{S,\psi}=\delta^{f_{1}}_{f_{3}}\delta^{f_{2}}_{f_{4}}+\delta^{f_{1}}_{f_{4}}\delta^{f_{2}}_{f_{3}},$ (12)
$T^{[f]}_{A,\psi}=\delta^{f_{1}}_{f_{3}}\delta^{f_{2}}_{f_{4}}-\delta^{f_{1}}_{f_{4}}\delta^{f_{2}}_{f_{3}},$ (13)
with $f_{1,2}$ the flavor indices for the field $\psi$ and $f_{3,4}$ for the field ${\psi^{\dagger}}$, and the subscript $S,A$ indicates that the tensor is symmetric or anti-symmetric under the permutation of 1,21,21 , 2 or 3,43,43 , 4. Similarly, the gauge tensor can also be classified as $T^{[g]}_{S}$ and $T^{[g]}_{A}$ as follows:
$T^{[g]}_{S,\psi}=\delta^{i_{1}}_{i_{3}}\delta^{i_{2}}_{i_{4}}+\delta^{i_{1}}_{i_{4}}\delta^{i_{2}}_{i_{3}},$ (14)
$T^{[g]}_{A,\psi}=\delta^{i_{1}}_{i_{3}}\delta^{i_{2}}_{i_{4}}-\delta^{i_{1}}_{i_{4}}\delta^{i_{2}}_{i_{3}},$ (15)
the $i$’s are gauge indices of the fields. For the Lorentz tensor there is only one anti-symmetirc contraction: $T^{[L]}_{A}=\epsilon^{\alpha_{1}\alpha_{2}}\tilde{\epsilon}^{\dot{\alpha}_{3}\dot{\alpha}_{4}}$ with $\alpha_{1,2}$ and $\dot{\alpha}_{3,4}$ being the spinor indices of $\psi$ and ${\psi^{\dagger}}$ respectively. In order to construct a totally anti-symmetric tensor regarding to permuting two $\psi$ or $\psi^{\dagger}$, $T^{[f]}_{S}$ ($T^{[f]}_{A}$) must be paired with $T^{[g]}_{S}$ ($T^{[g]}_{A}$), which results in two possible independent operators. To simplify our analysis we consider two independent linear combination of the operators, resulting in:
${\cal O}^{1}=({\psi^{\dagger}}^{f_{1}i_{1}}\bar{\sigma}^{\mu}\psi_{f_{1}i_{1}})({\psi^{\dagger}}^{f_{2}i_{2}}\bar{\sigma}^{\mu}\psi_{f_{2}i_{2}}),$ (16a)
${\cal O}^{2}=({\psi^{\dagger}}^{f_{1}i_{1}}\bar{\sigma}^{\mu}\psi_{f_{2}i_{1}})({\psi^{\dagger}}^{f_{2}i_{2}}\bar{\sigma}^{\mu}\psi_{f_{1}i_{2}}),$ (16b)
and the spinor indices in the parentheses are contracted implicitly. The exact same reasoning can be applied to the type ${\chi^{\dagger}}^{2}\chi^{2}$ as the Lorentz and flavor structures are the same as the type ${\psi^{\dagger}}^{2}\psi^{2}$. It can be proven that the gauge invariant tensors can be organized into two symmetric $T^{[g]}_{S,1/2,\chi}$ and one anti-symmetric $T^{[g]}_{A,\chi}$ for both antisymmetric and symmetric representations for $N_{c}\geq 4$. Therefore to form a totally anti-symmetric tensor, $T^{[g]}_{S,\{1,2\},\chi}$ should be paired with $T^{[f]}_{S\chi}$ and $T^{[g]}_{A,\chi}$ be paired with $T^{[f]}_{A,\chi}$, leading to three independent operators. On the other hand, $T^{[g]}_{S,\{1,2\},\chi}$ and $T^{[g]}_{A,\chi}$ can be explicitly constructed with $\delta$ and generator $T^{B}$ in the following form:
$T^{[g]}_{S,1,\chi}$ $=\delta^{a_{1}}_{a_{3}}\delta^{a_{2}}_{a_{4}}+\delta^{a_{2}}_{a_{3}}\delta^{a_{1}}_{a_{4}},$ (17)
$T^{[g]}_{S,2,\chi}$ $=T^{Ba_{1}}_{{\rm anti}\ a_{3}}T^{Ba_{2}}_{{\rm anti}\ a_{4}}+T^{Ba_{2}}_{{\rm anti}\ a_{3}}T^{Ba_{1}}_{{\rm anti}\ a_{4}},$ (18)
$T^{[g]}_{A,\chi}$ $=\delta^{a_{1}}_{a_{3}}\delta^{a_{2}}_{a_{4}}-\delta^{a_{2}}_{a_{3}}\delta^{a_{1}}_{a_{4}},$ (19)
where $a_{i}$ are the gauge indices for $\chi$ or ${\chi^{\dagger}}$, and repeated indices are contracted implicitly. From this equations, one can deduce that $\delta^{a_{1}}_{a_{3}}\delta^{a_{2}}_{a_{4}}$, $\delta^{a_{2}}_{a_{3}}\delta^{a_{1}}_{a_{4}}$ and $T^{Ba_{1}}_{a_{3}}T^{Ba_{2}}_{a_{4}}$ are also the three independent gauge tensors, which leads to the following three independent operators:
${\cal O}^{3}$ $=({\chi^{\dagger}}^{f_{1}a_{1}}\bar{\sigma}^{\mu}\chi_{f_{1}a_{2}})({\chi^{\dagger}}^{f_{2}a_{2}}\bar{\sigma}^{\mu}\chi_{f_{2}a_{1}}),$ (20a)
${\cal O}^{4}$ $=({\chi^{\dagger}}^{f_{1}a_{1}}\bar{\sigma}^{\mu}\chi_{f_{1}a_{1}})({\chi^{\dagger}}^{f_{2}a_{2}}\bar{\sigma}^{\mu}\chi_{f_{2}a_{2}}),$ (20b)
${\cal O}^{5}$ $=({\chi^{\dagger}}^{f_{1}a_{1}}\bar{\sigma}^{\mu}T^{Ba_{2}}_{a_{1}}\chi_{f_{1}a_{2}})({\chi^{\dagger}}^{f_{2}a_{3}}\bar{\sigma}^{\mu}T^{Ba_{4}}_{a_{3}}\chi_{f_{2}a_{4}}).$ (20c)
For the operator of the type ${\chi^{\dagger}}\chi{\psi^{\dagger}}\psi$, it is quite obvious that there are two ways to form gauge singlet – either with $\delta$’s or generators $T$ to contract two pairs of conjugate fields. Flavor-wise there is only one way to contract the indices, leading to 2 independent operators:
${\cal O}^{6}$ $=({\psi^{\dagger}}^{f_{1}i_{1}}\bar{\sigma}^{\mu}\psi_{f_{1}i_{1}})({\chi^{\dagger}}^{f_{2}a_{2}}\bar{\sigma}^{\mu}\chi_{f_{2}a_{2}}),$ (21a)
${\cal O}^{7}$ $=({\psi^{\dagger}}^{f_{1}i_{1}}\bar{\sigma}^{\mu}T^{Bi_{2}}_{i_{1}}\psi_{f_{1}i_{2}})({\chi^{\dagger}}^{f_{2}a_{3}}\bar{\sigma}^{\mu}T^{Ba_{4}}_{a_{3}}\chi_{f_{2}a_{4}}).$ (21b)
In the above notations, we write the generator of anti-symmetric representation as a matrix in ordinary way with two indices for the anti-symmetric representation. However, one can also use the notation of  7, where the field $\chi$ is written in terms of two anti-symmetric fundamental indices. In this notation, one can write the generator $T_{\rm anti}$ in terms of generator of fundamental representation and the $\delta$-function that contract a fundamental and an anti-fundamental indices. Formally, it can be expressed as follows:
$(T^{A}_{\rm anti})_{(i_{1},j_{1})}{}^{({i}_{2},{j}_{2})}=\delta^{{j_{2}}}_{j_{1}}T_{fi_{1}}^{A{i}_{2}}-\delta^{{i}_{1}}_{j_{1}}T_{fi_{1}}^{A{j}_{2}}-\delta^{{j}_{2}}_{i_{1}}T_{fj_{1}}^{A{i}_{2}}+\delta^{{i}_{2}}_{i_{1}}T_{fj_{1}}^{A{i}_{2}}\;,$ (22)
where $T_{f}$ is the generator of fundamental representation. The advantage to use this notation is that in the computation of the flow equation, the trace involving the antisymmetric generators can always be converted to the trace of the fundamental generators, which can be simplified with programs such as FeynCalc Shtabovenko et al. (2025).
Appendix B Full effective action and flow equations
[…]
$\partial_{t}\Gamma_{k}^{(\bar{\psi}_{a}\psi_{b}\bar{\psi}_{c}\psi_{d})}=-\partial_{t}\left(Z^{2}_{\psi}\,\lambda_{1}\right){\cal T}^{abcd}_{1,L}-\partial_{t}\left(Z^{2}_{\psi}\lambda_{2}\right){\cal T}_{2,L}^{abcd}\,,$ (23)
which results from taking four functional derivatives of the Wetterich equation with respect to fermionic fields. Here ${\cal T}^{\ldots}_{1/2}$ are the tensors structures constructed in (16) which contain gauge/flavor/Lorentz indices for the respective operators ${\cal O}_{1/2}$. To disentangle the dressing functions associated to the different tensor structures, one must employ appropriate projections to contract the open indices.
The most efficient strategy is to construct projection operators such that each projected correlator depends solely on one of the dressing functions Braun et al. (2025); Gehring et al. (2015). This is in fact, particularly relevant when bosonizing the theory. A simpler alternative, sufficient for the present purpose, is to construct a complete basis of projectors directly adapted to the tensor structures under consideration. We define:
${\cal P}_{i,R}^{abcd}={\cal T}_{i,R}^{badc}\,,$ (24)
as the projector which, when applied to (23), yield the projected flow equations:
$\partial_{t}\left[{\cal P}_{1,R}^{abcd}\Gamma_{k}^{(\bar{\psi}_{a}\psi_{b}\bar{\psi}_{c}\psi_{d})}\right]$ $=-\partial_{t}\left(Z^{2}_{\psi}\lambda_{1}\right)\left[{\cal P}_{1,R}^{abcd}{\cal T}_{1,L}^{abcd}\right]-\partial_{t}\left(Z^{2}_{\psi}\lambda_{2}\right)\left[{\cal P}_{1,R}^{abcd}{\cal T}_{2,L}^{abcd}\right]\,,$ (25a)
$\partial_{t}\left[{\cal P}_{2,R}^{abcd}\Gamma_{k}^{(\bar{\psi}_{a}\psi_{b}\bar{\psi}_{c}\psi_{d})}\right]$ $=-\partial_{t}\left(Z^{2}_{\psi}\lambda_{1}\right)\left[{\cal P}_{2,R}^{abcd}{\cal T}_{1,L}^{abcd}\right]-\partial_{t}\left(Z^{2}_{\psi}\lambda_{2}\right)\left[{\cal P}_{2,R}^{abcd}{\cal T}_{2,L}^{abcd}\right]\,.$ (25b)
The $R$ subindex stands for a replacement of $\bar{\sigma}^{\mu}\to\sigma^{\mu}$ in the respective chiral tensor structure. Solving (25) yields the flows of the dressing functions $\bar{\lambda}_{1}$ and $\bar{\lambda}_{2}$:
$\partial_{t}\bar{\lambda}_{1}$ $=(2+2\eta_{\psi})\bar{\lambda}_{1}+\frac{k^{2}}{Z_{\psi}^{2}}\frac{\partial_{t}\left[{\cal P}_{1,R}^{abcd}\Gamma_{k}^{(\bar{\psi}_{a}\psi_{b}\bar{\psi}_{c}\psi_{d})}\right](N_{c}N_{\psi}+1)-\partial_{t}\left[{\cal P}_{2,R}^{abcd}\Gamma_{k}^{(\bar{\psi}_{a}\psi_{b}\bar{\psi}_{c}\psi_{d})}\right](N_{c}+N_{\psi})}{128N_{c}(N_{c}^{2}-1)N_{\psi}(N_{\psi}^{2}-1)}\,,$ (26a)
$\partial_{t}\bar{\lambda}_{2}$ $=(2+2\eta_{\psi})\bar{\lambda}_{2}+\frac{k^{2}}{Z_{\psi}^{2}}\frac{\partial_{t}\left[{\cal P}_{2,R}^{abcd}\Gamma_{k}^{(\bar{\psi}_{a}\psi_{b}\bar{\psi}_{c}\psi_{d})}\right](N_{c}N_{\psi}+1)-\partial_{t}\left[{\cal P}_{1,R}^{abcd}\Gamma_{k}^{(\bar{\psi}_{a}\psi_{b}\bar{\psi}_{c}\psi_{d})}\right](N_{c}+N_{\psi})}{128N_{c}(N_{c}^{2}-1)N_{\psi}(N_{\psi}^{2}-1)}\,.$ (26b)
This procedure can be repeated for all other four-fermion operators to obtain their respective flows of the dressings.
The diagrammatic flows are computed in the vanishing external momentum limit, which, together with the choice of the Litim regulator grants us access to the following analytic expressions for the flows:
$\partial_{t}\bar{\lambda}_{1}=$ $(2+2\,\eta_{\psi})\bar{\lambda}_{1}+\frac{1}{16\pi^{2}}\bigg{[}\frac{g^{4}\left(3N_{c}^{2}+4\right)(5\eta_{A}+3\eta_{\psi}-45)}{160N_{c}^{2}}+\frac{g^{2}(5\eta_{A}+6\eta_{\psi}-60)(\bar{\lambda}_{1}-\bar{\lambda}_{2}N_{c})}{10N_{c}}$ (27a)
$-\frac{1}{20}(\eta_{\chi}-5)N_{\chi}\left(\bar{\lambda}_{6}^{2}(N_{c}-1)N_{c}+\frac{(N_{c}-2)}{2N_{c}}\bar{\lambda}_{7}^{2}\right)+\frac{4}{5}(\eta_{\psi}-5)\left(\bar{\lambda}_{2}^{2}-\bar{\lambda}_{1}\bar{\lambda}_{2}(N_{c}+N_{\psi})-\bar{\lambda}_{1}^{2}\frac{(N_{c}N_{\psi}-1)}{2}\right)\bigg{]}\,,$
$\partial_{t}\bar{\lambda}_{2}=$ $(2+2\,\eta_{\psi})\bar{\lambda}_{2}+\frac{1}{16\pi^{2}}\bigg{[}\frac{g^{4}\left(N_{c}^{2}-8\right)(5\eta_{A}+3\eta_{\psi}-45)}{160N_{c}}+\frac{g^{2}(5\eta_{A}+6\eta_{\psi}-60)(\bar{\lambda}_{2}-\bar{\lambda}_{1}N_{c})}{10N_{c}}$
$-\frac{1}{40}(\eta_{\chi}-5)\bar{\lambda}_{7}^{2}(N_{c}-2)N_{\chi}+\frac{2}{5}(\eta_{\psi}-5)\left(4\bar{\lambda}_{1}\bar{\lambda}_{2}-\bar{\lambda}_{2}^{2}(N_{c}+N_{\psi})\right)\bigg{]}\,,$ (27b)
$\partial_{t}\bar{\lambda}_{3}=$ $(2+2\,\eta_{\chi})\bar{\lambda}_{3}+\frac{1}{16\pi^{2}}\bigg{[}\frac{1}{10}g^{4}(5\eta_{A}+3\eta_{\chi}-45)+\frac{\left(g^{2}(5\eta_{A}+6\eta_{\chi}-60)\left(N_{c}\bar{\lambda}_{5}-2\bar{\lambda}_{3}\right)\right)}{5N_{c}}$
$+\frac{1}{5}(\eta_{\chi}-5)\left(8\bar{\lambda}_{4}\bar{\lambda}_{3}+\frac{4\left(N_{c}^{2}-N_{c}+4\right)}{N_{c}}\bar{\lambda}_{5}\bar{\lambda}_{3}-\bar{\lambda}_{3}^{2}\left(N_{c}^{2}-N_{c}+2N_{\chi}\right)+6\bar{\lambda}_{5}^{2}\right)\bigg{]}\,,$ (27c)
$\partial_{t}\bar{\lambda}_{4}=$ $(2+2\,\eta_{\chi})\bar{\lambda}_{4}+\frac{1}{16\pi^{2}}\bigg{[}\frac{g^{4}\left(N_{c}^{2}-2N_{c}-4\right)(5\eta_{A}+3\eta_{\chi}-45)}{10N_{c}^{2}}$
$-g^{2}\frac{\left(5\eta_{A}+6\eta_{\chi}-60\right)\left(\lambda_{5}\left(N^{2}{}_{c}-2N_{c}-4\right)+2\lambda_{3}N_{c}\right)}{5N_{c}^{2}}$ (27d)
$+\frac{\left(\eta_{\chi}-5\right)}{10}\left(\frac{48\bar{\lambda}_{5}\bar{\lambda}_{3}}{N_{c}}+\frac{\bar{\lambda}_{5}^{2}\left(12\left(N_{c}^{2}-2N_{c}-4\right)\right)}{N_{c}^{2}}+\frac{\bar{\lambda}_{4}\bar{\lambda}_{5}\left(8\left(-N_{c}^{2}+N_{c}+2\right)\right)}{N_{c}}\right)$
$+\frac{\left(\eta_{\chi}-5\right)}{10}\left(-4\bar{\lambda}_{4}\bar{\lambda}_{3}\left(N_{c}^{2}-N_{c}+2N_{\chi}\right)+\bar{\lambda}_{4}^{2}\left(-2N_{\chi}N_{c}^{2}+2N_{\chi}N_{c}+4\right)+8\bar{\lambda}_{3}^{2}\right)-\frac{\left(\eta_{\psi}-5\right)}{10}N_{\psi}N_{c}\bar{\lambda}_{6}^{2}\bigg{]}\,,$ (27e)
$\partial_{t}\bar{\lambda}_{5}=$ $(2+2\,\eta_{\chi})\bar{\lambda}_{5}+\frac{1}{16\pi^{2}}\bigg{[}\frac{g^{4}\left(N_{c}^{2}-8N_{c}-32\right)(5\eta_{A}+3\eta_{\chi}-45)}{80N_{c}}+\frac{g^{2}(5\eta_{A}+6\eta_{\chi}-60)(\bar{\lambda}_{5}(N_{c}+4)-N_{c}(\bar{\lambda}_{4}+\bar{\lambda}_{3}))}{5N_{c}}$
$+\frac{1}{5}(\eta_{\chi}-5)\left(8\bar{\lambda}_{4}\bar{\lambda}_{5}-4\bar{\lambda}_{3}\bar{\lambda}_{5}(N_{\chi}-3)-\frac{\left(N_{c}^{2}(N_{\chi}+3)-2N_{c}(N_{\chi}-1)+16\right)}{N_{c}}\bar{\lambda}_{5}^{2}\right)-\frac{1}{20}(\eta_{\psi}-5)\bar{\lambda}_{7}^{2}N_{\psi}\bigg{]}\,,$ (27f)
$\partial_{t}\bar{\lambda}_{6}=$ $(2+\,\eta_{\psi}+\,\eta_{\chi})\bar{\lambda}_{6}+\frac{1}{16\pi^{2}}\bigg{[}\frac{g^{4}\left(N_{c}^{2}-N_{c}-2\right)(10\eta_{A}+3(\eta_{\chi}+\eta_{\psi}-30))}{20N_{c}^{2}}$
$-\frac{g^{2}\bar{\lambda}_{7}\left(N_{c}^{2}-N_{c}-2\right)(5\eta_{A}+3(\eta_{\chi}+\eta_{\psi}-20))}{10N_{c}^{2}}+\frac{3(\eta_{\chi}+\eta_{\psi}-10)}{10}\left(\bar{\lambda}_{6}^{2}+\frac{\left(N_{c}^{2}-N_{c}-2\right)}{2N_{c}^{2}}\bar{\lambda}_{7}^{2}\right)$
$-\frac{1}{5}(\eta_{\chi}-5)\left(\bar{\lambda}_{4}\bar{\lambda}_{6}\left(N_{c}^{2}N_{\chi}-N_{c}N_{\chi}+2\right)+\bar{\lambda}_{3}\bar{\lambda}_{6}\left(N_{c}^{2}-N_{c}+2N_{\chi}\right)+\frac{2\left(N_{c}^{2}-N_{c}-2\right)}{N_{c}}\bar{\lambda}_{5}\bar{\lambda}_{6}\right)$
$-\frac{2}{5}(\eta_{\psi}-5)\left(\bar{\lambda}_{1}\bar{\lambda}_{6}(N_{c}N_{\psi}+1)+\bar{\lambda}_{2}\bar{\lambda}_{6}(N_{c}+N_{\psi})\right)\bigg{]}\,,$ (27g)
$\partial_{t}\bar{\lambda}_{7}=$ $(2+\,\eta_{\psi}+\,\eta_{\chi})\bar{\lambda}_{7}+\frac{1}{16\pi^{2}}\bigg{[}\frac{g^{4}\left(N_{c}^{2}-4N_{c}-16\right)(10\eta_{A}+3(\eta_{\chi}+\eta_{\psi}-30))}{80N_{c}}$
$+\frac{g^{2}(5\eta_{A}+3(\eta_{\chi}+\eta_{\psi}-20))(\bar{\lambda}_{7}(N_{c}+4)-2\bar{\lambda}_{6}N_{c})}{10N_{c}}$
$+\frac{(\eta_{\chi}+\eta_{\psi}-10)}{5}\left(3\bar{\lambda}_{6}\bar{\lambda}_{7}-\frac{\left(N_{c}^{2}+3N_{c}+12\right)}{4N_{c}}\bar{\lambda}_{7}^{2}\right)-\frac{2(\eta_{\psi}-5)}{5}\left(\bar{\lambda}_{1}\bar{\lambda}_{7}+\bar{\lambda}_{2}\bar{\lambda}_{7}N_{\psi}\right)$
$-\frac{2(\eta_{\chi}-5)}{5}\left(\bar{\lambda}_{4}\bar{\lambda}_{7}+\bar{\lambda}_{3}\bar{\lambda}_{7}N_{\chi}-\frac{\left(N_{c}^{2}(N_{\chi}+1)-2N_{c}(N_{\chi}+1)-4\right)}{10N_{c}}\bar{\lambda}_{5}\bar{\lambda}_{7}\right)\bigg{]}\,.$ (27h)
Here, $\eta_{A}$, $\eta_{\psi}$ and $\eta_{\chi}$ stand for the anomalous dimensions of the gauge and fermion fields respectively. These are computed from the respective two-point functions
$\eta_{A}=$ $-\frac{\partial_{t}Z_{A}}{Z_{A}}=-\left.\frac{\partial_{p^{2}}\,{\rm tr}\left[{\cal P}^{(AA)}\partial_{t}\Gamma^{(AA)}_{k}\right]}{Z_{A}\,{\rm tr}\left[{\left({\cal P}^{(AA)}\right)}^{2}\right]}\right|_{p=0}=\frac{g^{2}}{16\pi^{2}}\left[\left(\frac{31}{36}\eta_{A}-\frac{11}{3}\right)N_{c}-\frac{(\eta_{\chi}-4)}{6}(N_{c}-2)N_{\chi}-\frac{(\eta_{\psi}-4)}{6}N_{\psi}\right]\,,$ (28)
$\eta_{\psi}=$ $-\frac{\partial_{t}Z_{\psi}}{Z_{\psi}}=-\left.\frac{\partial_{p^{2}}\,{\rm tr}\left[{\cal P}^{(\psi\psi)}\partial_{t}\Gamma^{(\bar{\psi}\psi)}_{k}\right]}{Z_{\psi}\,{\rm tr}\left[{\left({\cal P}^{(\psi\psi)}\right)}^{2}\right]}\right|_{p=0}=\frac{g^{2}\left(N_{c}^{2}-1\right)}{32\pi^{2}N_{c}}\left[\frac{(\eta_{\psi}-4)}{4}-\frac{(\eta_{A}-5)}{5}\right]\,,$ (29)
$\eta_{\chi}=$ $-\frac{\partial_{t}Z_{\chi}}{Z_{\chi}}=-\left.\frac{\partial_{p^{2}}\,{\rm tr}\left[{\cal P}^{(\chi\chi)}\partial_{t}\Gamma^{(\bar{\chi}\chi)}_{k}\right]}{Z_{\chi}\,{\rm tr}\left[{\left({\cal P}^{(\chi\chi)}\right)}^{2}\right]}\right|_{p=0}=\frac{g^{2}\left(N_{c}^{2}-N_{c}-2\right)}{16\pi^{2}N_{c}}\left[\frac{(\eta_{\chi}-4)}{4}-\frac{(\eta_{A}-5)}{5}\right]\,,$ (30)
and here we resum them, which provides one of the sources for the non-perturbative character of this approach. The $\cal P^{(\ldots)}$ stand for the projectors onto the kinetic terms of the respective fields.
Furthermore, from the flows in (27), the previously discussed coefficients ($\boldsymbol{c}_{\ldots}$) can be read off as a function of the number of chiral matter. To facilitate the identification of each contribution, we have displayed the flows as a function of the fermion multiplicities given by the global symmetries,
$G_{f}=SU\left(N_{\psi}\right)\times SU\left(N_{\chi}\right)\times U\left(1\right)\;,$ (31)
where the anomaly-free combination fixes,
$N_{\psi}=N_{\textrm{gen}}\left(N_{c}-4\right)$ ${\rm and}$ $N_{\chi}=N_{\textrm{gen}}\,.$ (32)
So far, we have not specified the pure gauge and Dirac parts of the action. In this work we employ a truncation in this sectors which includes the interaction terms up to dimension four and the classical tensor structures which is the same one as in Goertz et al. (2024). Furthermore, in the gauge fixing sector we have used the Landau gauge. As an additional cross-check, we have derived the flow of the gauge couplings from all the included avatars and obtained the correct one-loop structure.
Figure 2: Value of the $\boldsymbol{c}_{{\rm A},i}$ coefficients which reflect the strength of the primer gauge corrections on the generation of the different four-fermion dressings. The sign crossing of $\boldsymbol{c}_{{\rm A},5}$ at $N_{c}\sim 11$ and $\boldsymbol{c}_{{\rm A},7}$ provides the identification of regimes and potential subdomains.
Some further remarks are in order. First, note that only $\boldsymbol{c}_{{\rm C},ijk}$ depends on $N_{\textrm{gen}}$, because the only source of $N_{\textrm{gen}}$ dependence in the flow arises from pure fermionic loops, which contribute exclusively to $\bar{\lambda}_{i}\bar{\lambda}_{j}$ terms.
Moreover, part of the system has a resemblance to QCD which permits to perform some cross checks. Operators ${\cal O}_{1}$ and ${\cal O}_{2}$ are equivalent to those known as (V-A) and (V-A)adj but here only with one chirality Gies and Jaeckel (2006); Braun (2012); Goertz et al. (2024). In fact, we find an agreement up to a factor 1/2 in the $N_{\psi}$ contributions in their respective flows, providing a consistency check to our computation.
Furthermore, this comparison also allows us to learn more details about the structure of gauge-fermion theories. The computation here presented also allows us to understand the transformation of QCD like results to the chiral limit. One may start with a Dirac QCD-like basis Gies and Jaeckel (2006); Braun (2012); Goertz et al. (2024) and decouple one of the chiralities. Note that this is merely useful to understand the structural level as the product theories are anomalous. The first consequence is that the scalar-pseudoscalar and vector+axial channels are not generated at all as required the presence of two chiralities, as the diagrams are all proportional to $N_{L}N_{R}$. As seen here, in the QCD-like scenario the ${\cal O}_{({\rm V-A})}$ and ${\cal T}_{({\rm V-A})_{{\rm adj}}}$ channels (analogues to our ${\cal O}_{1}$ and ${\cal T}_{2}$) do not fulfill the lowest-order structure for triggering a resonance. Nonetheless, as discussed in Goertz et al. (2024), in the many-flavor QCD limit with locking dynamics evidences for a further resonant ${\cal T}_{({\rm V-A})_{{\rm adj}}}$ channels and condensates has been found, mainly led by strong fermion dynamics in the (S-P) channel. Moreover, the remaining vector-axial structure is indeed generated leading to the results here obtained. A last remark is in order, in the presence of one single chirality the charge-parity symmetry is not present and terms that would violate it in vector-like theories can appear. For example, the four-fermion interactions now receive contributions of the form
$(N_{L}-N_{R})\,\,\epsilon^{\mu\nu\sigma\rho}p_{1}^{\mu}p_{2}^{\nu}p_{3}^{\sigma}q^{\rho}$ (33)
which in the QCD limit exactly vanish but here preserve. The $p_{i}$’s stand for the external momenta of the four-point function which, in the present vanishing external-momenta approximation such contributions drop. Nonetheless, we highlight this type of novel higher-order corrections encoded in the momentum dependence of the flows.
Appendix C A closer look at particular examples
To better illustrate the underlying dynamics and the analysis performed in this work, this appendix presents a detailed discussion of the integrated RG flows for selected representative theories. We simultaneously integrate the flows in (27) along with the flow of the gauge coupling. The four-fermion flows incorporate the resumed form of the anomalous dimensions, which effectively accounts for infinite-order corrections.
As previously discussed, we derived the RG flow for various avatars of the gauge coupling and confirmed that the one-loop structure is the correct one. Nevertheless, for the purposes of this appendix, we employ the multi-loop perturbative RG equations, which will be presented in Appendix D. This choice implies that the four-fermion couplings do not feed back self-consistently into the evolution of the gauge coupling. However, the equivalent higher-order effects are already encoded in the perturbative running. While performing the fully self-consistent integration poses no technical difficulty, for the qualitative discussion at hand this simplification introduces no significant differences. In fact, using the perturbative expression allows us to correctly implement the multi-loop symmetry factors, which are relevant for fixed-point determination and for connecting to the conformal regime.
We emphasize that the analysis in the main text does not rely on this approximation. The determination of $\alpha^{\textrm{crit}}_{\textrm{SB}}$ depends solely on the flow of the four-fermion dressings, which are only affected by the truncation through the gauge-field anomalous dimension. Note that we are not attempting to capture the full IR dynamics, and in particular we have neglected confinement and the formation of a mass gap in the gauge sector.
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
Figure 3: Integrated RG flows for different theories, specified by various values of $N_{c}$ and $N_{\textrm{gen}}$. Top panels show the evolution of the gauge coupling; bottom panels show the four-fermion couplings defined in (6). Insets display the absolute value of the four-fermion couplings on a logarithmic scale.
Having set the stage, we now integrate the coupled system of four-fermion and gauge flows for different values of $N_{c}$ and $N_{\textrm{gen}}$, as shown in Figure 3. As detailed in the main text, the emergence of an IR fixed point in the flow of the four-fermion couplings indicates that these are fully determined by gauge dynamics. Furthermore, setting different UV boundary conditions does not affect the IR behavior: the system quickly approaches a partial fixed point, as observed in related contexts Goertz et al. (2024); Braun et al. (2016). However, if the UV boundary conditions are too large, e.g., $\bar{\lambda}_{i,\,{k\to\infty}}\sim\mathcal{O}(10)$, the theory exits the universality class of perturbatively renormalizable gauge-fermion systems and enters the domain of nonperturbatively non-renormalizable gauge-NJL models. In that case, the IR behavior depends on the UV data, reflecting the effective nature of these theories. To avoid such effects, we take $\bar{\lambda}_{i,\,{k\to\infty}}=0$ and $\alpha_{g,\,{k\to\infty}}\simeq 0$.
As the theory flows away from the UV Gaussian fixed point, an infinite tower of operators is dynamically generated, which is reflected in the growth of their associated couplings. The precise behavior depends on the symmetry structure and corresponding quantum fluctuations, as illustrated in Figure 3.
Due to the self-interaction terms $\boldsymbol{c}_{{\rm C},iii}$, the flow equations take a parabolic form (see Figures 4 and 5), where the direction of the parabola depends on the sign of the coefficient. Starting from the UV, the flow for small $\alpha_{g}$ always admits an IR attractive fixed point, which continuously connects to $\bar{\lambda}_{i,\,{k\to\infty}}=0$ as $\alpha_{g}\to 0$. This behavior is insensitive to the coefficients $\boldsymbol{c}_{{\rm A}},\boldsymbol{c}_{{\rm B}},\boldsymbol{c}_{{\rm C}}$, provided $\alpha_{g}$ remains sufficiently small.
However, if any of the couplings displays a resonant structure (see Section IV) and $\alpha_{g}$ reaches $\alpha^{\textrm{crit}}_{\textrm{SB}}$, the corresponding IR fixed point disappears via a merger and a singularity appears in the RG flow. As shown in the main text, for $N_{c}=5$, only two channels meet this condition: ${\cal O}_{5}$ and ${\cal O}_{7}$.
More precisely, $\alpha^{\textrm{crit}}_{\textrm{SB}}$ is reached first, and the singularity appears at a slightly lower scale, signaling the onset of dynamical symmetry breaking at the scale $k_{\textrm{dSB}}$, see Goertz et al. (2024); Ihssen et al. (2024). Note that the scale where $\alpha^{\textrm{crit}}_{\textrm{SB}}$ is reached and $k_{\textrm{dSB}}$ are not identical. The gap between them can grow if other competing dynamics, such as confinement, become important, particularly when the system lies far from the conformal window.
Moving across theory space, the evolution of the gauge coupling changes, as fermionic contributions decelerate its IR growth. Eventually, as $N_{\textrm{gen}}$ increases, the gauge coupling flows to an IR fixed point that remains below $\alpha^{\textrm{crit}}_{\textrm{SB}}$, and no dynamical scale is generated. This scenario is shown in the right-most panel of Figure 3, where all four-fermion couplings approach scale-invariant values as $k\to 0$. Conversely, slightly decreasing $N_{\textrm{gen}}$ may place the IR fixed point just above $\alpha^{\textrm{crit}}_{\textrm{SB}}$, leading to walking behavior and eventual condensation at $k_{\textrm{dSB}}$, as illustrated in the central panel of Figure 3 and discussed in Goertz et al. (2024); Miransky and Yamawaki (1997); Miransky (1999); Braun and Gies (2006, 2010); Braun et al. (2011). Furthermore, as $k_{\textrm{dSB}}$ is approached, all of the couplings start rapidly growing, see left and middle plots in Figure 3. This is an artifact of the four-fermion computation. When bosonising, even partially only the dominant channel, it is known that the subleading tensor structures are not driven to a singularity, see e.g. Goertz et al. (2024); Jaeckel and Wetterich (2003); Denz et al. (2020); Ihssen et al. (2024).
To better visualize the dynamical symmetry breaking, we show in Figures 4 and 5 the beta functions of the four-fermion couplings plotted against themselves for the case $N_{c}=5$, $N_{\textrm{gen}}=7$ (central panel of Figure 3). The flows are displayed at different RG times, from UV (blue) to IR (red), terminating at $k=k_{\textrm{dSB}}$. Here we see that the fixed-point merger, which as we discussed, is predominantly driven by the lowest-order gauge corrections and to a lesser extent by higher-order corrections via other channels.
By inspecting these flows, it is straightforward to identify the resonant channels. For example, in the cases of $\bar{\lambda}_{5}$ and $\bar{\lambda}_{7}$, the fixed point disappears in the IR, whereas for other couplings the fixed point persists. As a result, no condensation is expected in channels like ${\cal O}_{1}$ and ${\cal O}_{2}$, which only involve $\psi$ fields.
Figure 4: Flow of the $\bar{\lambda}_{5}$ dressing as a function of itself evaluated at different RG times for a theory with $N_{c}=5$ and $N_{\textrm{gen}}=7$. Different curves and points indicate the running and position of the coupling at different scales.
Figure 5: Same as Figure 4, but for the remaining four-fermion dressings defined in (6).
Finally, we comment on the signs of the four-fermion couplings defined in (5). These are dictated by the lowest-order contributions shown in Figures 2 and 9. For $N_{c}=5$, all couplings except $\bar{\lambda}_{5}$ and $\bar{\lambda}_{7}$ have positive signs in our convention. This choice is motivated by bosonization arguments, where positive sign dressings correspond to a positive curvature of the mesonic potential in the chirally symmetric regime. These non-resonant channels reflect precisely this behavior. Conversely, the resonant channels exhibit negative sign dressings, which in a naive bosonization scheme would imply a tachyonic instability. However, this is not a concern here, as bosonization itself can introduce additional negative signs. Moreover, similar behavior is observed in QCD, where negative-sign couplings coexist with well-understood IR physics.
Appendix D Gauge beta functions and infrared fixed points
The $\overline{\mathrm{MS}}$ beta function for a $SU(N_{c})$ chiral gauge theory can be written as
$\beta(g)=$ $-g\,\sum_{n=1}^{\infty}\left(\frac{g^{2}}{16\pi^{2}}\right)^{n}\,\beta_{n}\,,$ (34)
with the first four-loop coefficients given by
$\beta_{0}$ $=\frac{11}{3}\,C_{2}(G)-\frac{2}{3}\sum_{i}T(R_{i})N_{R_{i}}\,,$ (35a)
$\beta_{1}$ $=\frac{34}{3}\,C_{2}(G)^{2}-\sum_{i}\left[2\,C_{2}(R_{i})+\frac{10}{3}\,C_{2}(G)\right]T(R_{i})N_{R_{i}}$ (35b)
$\beta_{2}$ $=\frac{2857}{54}\,C_{2}(G)^{3}-\sum_{i}\left[-C_{2}(R_{i})^{2}+\frac{205}{18}\,C_{2}(G)C_{2}(R_{i})+\frac{1415}{6}\,C_{2}(G)^{3}\right]T(R_{i})N_{R_{i}}$
$\;\;\;\;\;+\sum_{i,j}\left[\frac{11}{9}\,C_{2}(R_{i})+\frac{79}{54}\,C_{2}(G)\right]T(R_{i})T(R_{j})N_{R_{i}}N_{R_{j}}$ (35c)
$\beta_{3}$ $=\left(\frac{150653}{486}-\frac{44}{9}\zeta_{3}\right)C_{2}(G)^{4}-\left(\frac{80}{9}-\frac{704}{3}\zeta_{3}\right)\frac{d_{AA}}{{N_{c}}^{2}-1}$
$\;\;\;\;\;-\sum_{i}\left[-23C_{2}(R_{i})^{3}+\left(\frac{4204}{54}-\frac{352}{18}\zeta_{3}\right)C_{2}(G)C_{2}(R_{i})^{2}-\left(\frac{7073}{486}-\frac{656}{18}\zeta_{3}\right)C_{2}(G)^{2}C_{2}(R_{i})\right.$
$\;\;\;\;\;+\left.\left(\frac{39143}{162}-\frac{136}{6}\zeta_{3}\right)\,C_{2}(G)^{3}\right]T(R_{i})N_{R_{i}}+\sum_{i}N_{R_{i}}\left(\frac{512}{18}-\frac{1664}{6}\zeta_{3}\right)\frac{d_{R_{i}A}}{{N_{c}}^{2}-1}$
$\;\;\;\;\;-\sum_{i,j}\left[-\left(\frac{184}{12}-\frac{64}{4}\zeta_{3}\right)C_{2}(R_{i})C_{2}(R_{j})+\left(\frac{304}{108}+\frac{128}{36}\zeta_{3}\right)C_{2}(R_{i})^{2}-\left(\frac{17152}{972}+\frac{448}{36}\zeta_{3}\right)C_{2}(R_{i})C_{2}(G)\right.$
$\;\;\;\;\;-\left.\left(\frac{7930}{324}+\frac{224}{36}\zeta_{3}\right)C_{2}(G)^{2}\right]T(R_{i})T(R_{j})N_{R_{i}}N_{R_{j}}-\sum_{i,j}N_{R_{i}}N_{R_{j}}\left(\frac{704}{36}-\frac{512}{12}\zeta_{3}\right)\frac{d_{R_{i}R_{j}}}{{N_{c}}^{2}-1}$
$\;\;\;\;\;+\sum_{i,j,k}N_{R_{i}}N_{R_{j}}N_{R_{k}}T(R_{i})T(R_{j})T(R_{k})\left(\frac{154}{243}C_{2}(R_{i})\frac{53}{243}C_{2}(G)\right)\,.$ (35d)
[…]
Representation Dimension Quadratic Casimir $C_{2}(R)$ Dynkin Index $T(R)$
Anti-Fundamental $N$ $\frac{N^{2}-1}{2N}$ $\frac{1}{2}$
Adjoint $N^{2}-1$ $N$ $N$
Antisymmetric $\frac{N(N-1)}{2}$ $\frac{(N-2)(N+1)}{N}$ $\frac{N-2}{2}$
Symmetric $\frac{N(N+1)}{2}$ $\frac{(N+2)(N-1)}{N}$ $\frac{N+2}{2}$
Table 2: Dimensions, quadratic Casimirs and Dynkin indices for various representations of $\mathrm{SU}(N)$. Conjugate representations have identical values.
  Anti-Fund   Adjoint Antisymmetric
Anti-Fund $\frac{({N}^{6}-7{N}^{4}+24{N}^{2}-18)}{96{N}^{2}}$ $\frac{{N}({N}^{4}+5{N}^{2}-6)}{48}$ $\frac{({N}-2)({N}^{2}-1)({N}(N-6)^{2}(N+6)+72)}{96{N}^{2}}$
Adjoint $\frac{{N}^{2}({N}^{4}+35{N}^{2}-36)}{24}$ $\frac{({N}-2)({N}^{2}-1){N}((N-6)N+24)}{48}$
  Antisymmetric   $\frac{(N-2)(N-1)(N+1)((N-2)N(N((N-12)N+48)+144)-576)}{96N^{2}}$
Table 3: Group coefficient $d_{R_{f}R_{f^{\prime}}}$ for different $\mathrm{SU}(N)$ representations.